A quality-control manager at an amusement park feels that the amount of time that people spend waiting in line to ride the roller coaster is too long. To determine if a new loading and unloading procedure is effective in reducing the wait time, she measures the amount of time (in minutes) people are waiting in line for seven days. To make a reasonable comparison, she chooses times when whether conditions are similar.

        

                                                 Mon       Tues      Wed      Thurs       Fri         Sat          Sun

                                             2 p.m.    2 p.m.    2 p.m.    2 p.m.    2 p.m.    4 p.m.    12 noon


          Wait time before, 𝑋1       12          26          20          38          57          82            57

          Wait time after, 𝑋2          11          28          19          36          59          75            55

         

 

Is the new loading and unloading procedure effective in reducing the wait time at the α = 0.05 level of significance?

 

 

a.     State the hypothesis and identify the claim of the researcher.

b.     Find the critical value(s).

c.     Compute the test value.

d.     Make a decision on the null hypothesis. 

e.     Make a decision on the claim of the researcher. 


1
Expert's answer
2021-12-22T13:07:10-0500

Paired sample T-test

a.

"H_0:\mu_1=\mu_2" , new loading and unloading procedure is not effective in reducing the wait time (average amounts of wait time are equal)

"H_a:\mu_2<\mu_1" , new loading and unloading procedure is effective in reducing the wait time (average amount of wait time "after" is less than average amount of wait time "before")


b.

"df=n-1=6"

critical value:

"t_{crit}=2.232"


c.

test value:

"t=\frac{\mu_2-\mu_1}{s_d/\sqrt n}=\frac{-1.29}{1.149/\sqrt 7}=-1.119"

where sd is standard deviation of the difference


d.

Since "|t|<t_{crit}" we accept null hypothesis


e.

new loading and unloading procedure is not effective in reducing the wait time